OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

Let n = the largest number Let n-1= second larrgest, consecutive Let n -2 = smallest of those consecutive #s n + (n-1) + n(-2) = 267 n+ n + n -3 = 267 3n = 270 n= ? Can you find it now?

OpenStudy (anonymous):

Maybe I'm not sure yet but thanks!

OpenStudy (anonymous):

+ (n-2) **

OpenStudy (anonymous):

Is there anything confusing? Lemme know.

OpenStudy (anonymous):

n= 90. so now you can check if you're unsure. biggest #= 90 next biggest = 89 smallest = 88 90 + 88 + 89 = 267 checked

OpenStudy (anonymous):

Thanks so much do you mind explaining the steps? I have a test on this in a week.

OpenStudy (anonymous):

Ok , you want find a number right? Let that number be a n. Is this clear?

OpenStudy (anonymous):

"Now it says the sum of three consectutive numbers." sum means adding them. consecutive means that the number are next to each other.

OpenStudy (anonymous):

so, let n be your biggest number.

OpenStudy (anonymous):

got it so far

OpenStudy (anonymous):

the number that's consecutive (next to) and smaller than n is n-1. LIke the consecutive, smaller number of 10 is 9 (for example.) so we have n and n-1 so far.

OpenStudy (anonymous):

okay i think im getting it thanks!

OpenStudy (anonymous):

and the number that's consecutive and smaller than n-1, is n-2. so we have n , n-1, and n-2 now to find the sum , we add these 3 together.

OpenStudy (anonymous):

n + n-1 + n-2 = 267 and from there you solve the problem.

OpenStudy (anonymous):

@rational are you on stalking mode? xDD (I just noticed)

OpenStudy (rational):

lol just entered this thread as the problem seems interesting, want me leave ha :o

OpenStudy (anonymous):

oh sorry I didn't mean it that way! xDD Come back to me, rational!

OpenStudy (rational):

jk haha! it seems sum of 3 consecutive integers is always odd hmm

OpenStudy (rational):

another (similar) way @AndreaEwen (n-1) + (n) + (n+1) = 267 3n = 267 n = 89 so the required numbers are : 88,89,90 you can pick the largest number^^

OpenStudy (anonymous):

thank you! @rational

OpenStudy (anonymous):

good thinking! :)